# microtypeUnable to apply patch

### v1

@@ Abstract
Empathetic response generation is a desirable aspect of conversational agents, crucial for facilitating engaging and emotionally intelligent multi-turn conversations between humans and machines. Leveraging large language models for this task has shown promising results, yet challenges persist in ensuring both the empathetic quality of the responses and retention of the generalization performance of the models. In this paper, we propose a novel approach where we construct theory-driven preference datasets and use them to align LLMs with preference optimization algorithms to address these challenges. To measure empathetic response generation, we employ the EmpatheticDialogues dataset, assessing empathy with the diff-Epitome and BERTscore metrics, and evaluate the generalization performance on the MMLU benchmark. We make all datasets, source code, and models publicly available.111github.com/ondrejsotolar/empo
@@ 1 Introduction
Empathetic response generation (ERG) focuses on tuning a conversational agent toward understanding the user’s situation, feelings, and experience to generate appropriate, human-like responses. This goal was first attempted by the earliest rule-based chatbots like ELIZA (Weizenbaum, 1966), then more recently with deep learning approaches such as BlenderBot (Roller et al., 2021; Xu et al., 2022; Komeili et al., 2022; Shuster et al., 2022). The introduction of instruction-tuned large language models (LLMs) such as ChatGPT OpenAI (2023) has challenged the contemporary approaches to ERG. These models can participate in multi-turn conversations without additional training and are implicitly capable of generating empathetic conversations (Lee et al., 2022a).
@@
LLMs also have shown tremendous generalization capabilities and can draw on world knowledge obtained by training on Internet-size data. These traits were not present in the earlier non-LLM conversational agents. Recent developments focus on approaches that keep the generalization abilities of LLMs while improving empathy in responses, including prompt-engineering methods that do not change the model’s weights but rather tune the prompt content to assist the generation by in-context learning (Wang et al., 2024), often assisted by retrieval (Qian et al., 2023).
@@
In the current study, we propose a data-driven solution for ERG with LLMs, based on preference optimization: aligning LLMs via preference optimization algorithms. First, we build a preference dataset using the benchmark EmpatheticDialogues dataset (Rashkin et al., 2019). The dataset contains short multi-turn human-to-human dialogues grounded by emotion labels. We leverage this property to extract responses from the corpus of the polar opposite emotion label using Plutchik’s wheel (Plutchik, 2001) such that each prompt is paired with preferred and non-preferred completions. We then fine-tune a foundational LLM using Direct Preference Optimization (Rafailov et al., 2024) to generate responses aligned with the preferred candidate response.
@@
Using diff-Epitome (Lee et al., 2022b), a model-based empathy metric for multi-turn conversations, we show that training LLMs with our preference dataset improves ERG. We also investigate the models’ general language understanding using the MMLU benchmark to show how our method impacts the performance on other tasks. Our method is analogous to providing guardrails with helpful/harmful preference datasets (Bai et al., 2022). Contemporary prompt-engineering methods or additional training can be applied to models aligned in this way to adapt them further for any task. We also share novel observations from searching over the hyperparameter configuration space and provide code to apply our method to other datasets and models.
@@ 2 Related Work
Systems specialized for ERG such as KEMP (Li et al., 2022), CEM (Sahand Sabour, 2021), MIME (Majumder et al., 2020), EmpDG (Li et al., 2020), and DIFFUSEMP (Bi et al., 2023) or universal conversational agent like Blenderbot were not built for general instruction-following. Recent approaches such as (Wang et al., 2024; Qian et al., 2023; Li et al., 2024) employ prompt engineering with LLMs to some success. However, their approach differs principally from ours; we restructure the problem of ERG as an LLM alignment problem and produce a method for creating preference datasets and a modeling approach rather than a prompt construction algorithm.
@@
Several dialogue datasets, including IEMOCAP (Busso et al., 2008), MELD (Poria et al., 2019), DailyDialog (Li et al., 2017), Herzig et al., 2016, EmotionLines (Hsu et al., 2018), EmpathicReactions (Buechel et al., 2018), and EmoContext (Chatterjee et al., 2019), contain emotion labels. However, these datasets are either labeled with only a small set of emotions, limited by size, or lack a multi-turn character.
The EmpatheticDialogues (ED) is a dataset with 25K human-to-human dialogues and 32 emotion labels, which are derived from biological responses (Ekman, 1992; Plutchik, 1980) to larger sets of subtle emotions derived from contextual situations (Skerry and Saxe, 2015). It is the benchmark dataset on empathetic conversation as it addresses preceding datasets’ limitations.
Earlier deep learning-based conversational agents required large datasets for training from scratch, often at the expense of quality. A subset of ED was analyzed for listener-specific response intents, creating the EDOS dataset (Welivita et al., 2021), which includes 1M movie dialogues annotated with a BERT-based classifier. Recently, synthetic datasets like Chinese SMILE (Qiu et al., 2024) and Korean SoulChat (Chen et al., 2023) have been generated using LLMs.
@@
Evaluating ERG
Evaluation of empathy in dialogues has primarily focused on human evaluation and lexical overlap/semantic similarity with reference dialogues. The former is limited by the highly subjective nature of the task, the need to train annotators, and, usually, small sample sizes. We have found the latter to have little relevance for measuring empathy, given the highly creative nature of generative LLM responses, in agreement with the findings of Liu et al., 2016. We instead rely on model-based empathy metrics, which use classification models trained specifically for the task of measuring empathy in multi-turn conversations: diff-EPITOME (Lee et al., 2022b) which is based on the EPITOME (Sharma et al., 2020) classifiers. For measuring the generalization abilities, we use the standard benchmarks such as MMLU (Hendrycks et al., 2020), using the lm-evaluation-harness (Bommasani et al., 2023).
@@ 3 Experimental Setup
We conduct our experiments using the Zephyr-7B (Tunstall et al., 2023) model, a variant of the Mistral-7B foundational LLM (Jiang et al., 2023) fine-tuned for multi-turn dialogues using the UltraChat dataset (Ding et al., 2023). The smallest variant with 7 billion parameters already provides good general language understanding ability as it matches LLama-2-70B on many NLP benchmarks (Beeching et al., 2023). For all training steps, we explore the hyperparameter configuration space and optimize for a minimal impact on generalization while improving empathy scores.
@@
As Tunstall et al., 2023 showed, SFT is a necessary first step in alignment, which ensures that the preference dataset is in-domain for the aligned model. We perform SFT using the standard causal auto-regressive objective using the individual dialogues from EmpatheticDialogues. As per the definition of the dataset, the odd turns are considered the "user prompts" and the even turns the "responses". We limit the SFT training to the even-indexed turns by masking the odd turns to ignore them in the loss-function computation. For computational efficiency, we fine-tune using LoRA adapters (Hu et al., 2021).
@@
For constructing the completion pairs, we leverage a property of the ED dataset: each dialogue is associated with an emotion label. These were used as grounding during the ideation of each dialogue. For the preferred completion, we take the ground truth - the original response. For its rejected counterpart, we use Plutchik’s wheel of emotions (Figure 4), and the derivative emotional dyads (Figure  5,  Plutchik, 2001) to find the polar opposite emotion labels resulting in a lookup table (see A.3). For each completion, we randomly select one from the group of completions labeled with this opposite label. Because the stability can suffer from this random selection, we draw a fresh random completion for each new epoch of training and search for the hyperparameter configuration to offset the repetition of the preferred completions.
Model MMLU (5s) $\uparrow$ diff-ER $\downarrow$ diff-EX $\downarrow$ diff-IP $\downarrow$ FBert $\uparrow$
@@
Table 1: Language understanding, empathy, and semantic similarity with EmpatheticDialogues. All presented results are means ($\pm$ SD) of scores over multiple (4) training runs with the same hyperparameters but different seeds.
Empathy Evaluation
For measuring empathy, we use the diff-Epitome metric Lee et al. (2022a): a model-based metric specifically developed for measuring empathy in dialogues. diff-Epitome is an evolution of the Epitome classifiers (Sharma et al., 2020), which measure empathy in dialogue on a scale (0-2) from none (0), through weak (1), to strong (2) on three dimensions: empathetic responses (ER), explanations (EX), and interpretations (IP). The diff-Epitome uses a similar but open-ended continuous scale (0-) on the same dimensions by averaging the scores across the entire dataset, thus providing a measure of difference from the ground truth. Given our data-driven approach, we consider a lower difference from the ground truth better.
In preliminary human-evaluation experiments with lexical-overlap metrics such as BLEU or ROUGE and vector-similarity such as BERTscore, we found them uncorrelated with the perceived quality of the generated responses. This was especially evident with the largest LLMs, such as GPT4 and Claude, which generated high-quality responses yet received low scores from overlap metrics. We attribute this to the highly diverse nature of LLM generations. However, we keep using the semantic similarity BERTscore for sanity and increased metric diversity.
@@
We measure the general language of the models with established benchmarks, such as the Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2020) MMLU is designed to assess how LLMs learn and apply knowledge across a variety of domains by measuring their performance on 57 different multiple-choice style question tasks from STEM, humanities, and world knowledge. This benchmark requires the model to be capable of general instruction following.

### v2

@@ Abstract
Empathetic response generation is a desirable aspect of conversational agents, crucial for facilitating engaging and emotionally intelligent multi-turn conversations between humans and machines. Leveraging large language models for this task has shown promising results, yet challenges persist in ensuring both the empathetic quality of the responses and retention of the generalization performance of the models. We propose a novel approach where we construct theory-driven preference datasets based on emotion grounding and use them to align LLMs with preference optimization algorithms to address these challenges. To evaluate empathetic response generation, we employ the EmpatheticDialogues dataset, assessing empathy with the diff-Epitome and BERTscore metrics and with multi-dimensional human evaluation. Additionally, we measure diversity and emotional valence using feature-based methods. We also evaluate the impact of training on the generalization performance using the MMLU benchmark and tasks from the Open LLM Leaderboard. The results show that LLMs can be aligned for empathetic response generation by preference optimization while retaining their general performance and that emotion grounding can guide preference dataset creation.
@@ 1 Introduction
Empathetic response generation (ERG) focuses on tuning a conversational agent toward understanding the user’s situation, feelings, and experience to generate appropriate, human-like responses. This goal was first attempted by the earliest rule-based chatbots like ELIZA (Weizenbaum, 1966), then more recently with deep learning approaches such as used by the chatbot BlenderBot (Roller et al., 2021; Xu et al., 2022; Komeili et al., 2022; Shuster et al., 2022). The introduction of instruction-tuned large language models (LLMs) such as ChatGPT OpenAI (2023) has challenged the contemporary approaches to ERG. These models can participate in multi-turn conversations without additional training and are implicitly capable of generating empathetic conversations (Lee et al., 2022a).
@@
LLMs also have shown tremendous generalization capabilities and can draw on world knowledge obtained by training on Internet-size data. These traits were not present in the earlier non-LLM conversational agents. Recent developments focus on approaches that retain the generalization abilities of LLMs while improving empathy in responses, including prompt-engineering methods that do not change the model’s weights but rather tune the prompt content to assist the generation by in-context learning (Wang et al., 2024a), often assisted by retrieval (Qian et al., 2023).
@@
In the current study, we propose a data-driven solution for ERG with LLMs: aligning LLMs via preference optimization algorithms. First, we build a preference dataset using the benchmark dataset EmpatheticDialogues (Rashkin et al., 2019). It contains short multi-turn human-to-human dialogues grounded by emotion labels. We leverage this emotion grounding to sample dialog completions labeled with polar opposite emotions using Plutchik’s wheel (Plutchik, 2001) such that each prompt is paired with preferred and non-preferred completions. We then fine-tune a foundational LLM using Direct Preference Optimization (DPO) (Rafailov et al., 2024) to generate responses aligned with the preferred candidate response.
@@
Using metrics such as the model-based empathy metric for multi-turn conversations diff-Epitome (Lee et al., 2022b), several feature-based metrics for diversity and emotional valence, and multi-dimensional human evaluation, we show that training LLMs with our preference dataset helps to generate diverse and empathetic responses. We also investigate the models’ general language understanding using benchmarks such as the Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2020) and the tasks from the Open LLM Leaderboard (Beeching et al., 2023a). By monitoring how the training impacts language understanding performance, we show that it is possible to find hyperparameters that align the model for ERG while preventing catastrophic forgetting. Our method is analogous to providing guardrails with helpful/harmful preference datasets (Bai et al., 2022). Models aligned in this way can be further adapted for any task by contemporary prompt-engineering methods or additional training which is an advantage over alignment with methods that do not train the model weights. We also share novel observations from searching over the hyperparameter configuration space and provide code to apply our method to other datasets and models. We make all datasets, source code, and models publicly available.111github.com/justtherightsize/empo
@@ 2 Related Work
Systems specialized for ERG such as KEMP (Li et al., 2022), CEM (Sahand Sabour, 2021), MIME (Majumder et al., 2020a), EmpDG (Li et al., 2020), and DiffusEmp (Bi et al., 2023) or universal conversational agents like BlenderBot were not built for general language understanding or instruction-following. They can be evaluated on dialog completion, but they fail on other tasks as opposed to LLMs; thus, they are not solutions to the same problem.
Recent approaches to ERG such as Wang et al. (2024a); Qian et al. (2023); Li et al. (2024) employ LLMs enhanced by prompt engineering with some success. However, their approach differs principally from ours; we restructure the problem of ERG as an LLM alignment problem and produce a method for creating preference datasets and a modeling approach rather than a prompt construction algorithm.
@@
Several dialogue datasets, including IEMOCAP (Busso et al., 2008), MELD (Poria et al., 2019), DailyDialog (Li et al., 2017), Herzig et al., 2016, EmotionLines (Hsu et al., 2018), EmpathicReactions (Buechel et al., 2018), EmoContext (Chatterjee et al., 2019), and ESConv (Liu et al., 2021) contain emotion labels. However, these datasets are either labeled with only a small set of emotions, limited by size, or lacking a multi-turn character.
EmpatheticDialogues (ED) is a dataset with 25K human-to-human dialogues and 32 emotion labels, which are derived from biological responses (Ekman, 1992; Plutchik, 1980) to larger sets of subtle emotions derived from contextual situations (Skerry and Saxe, 2015). It has become the benchmark dataset on empathetic conversation because it addresses the limitations of the preceding datasets.
Earlier deep learning-based conversational agents required large datasets for training from scratch to learn to participate in dialogue, often at the expense of the dataset quality. An example is the EDOS dataset (Welivita et al., 2021) of 1M movie dialogues with emotion grounding annotated by classifiers trained on ED.
@@
Recently, it became feasible to generate high-quality synthetic datasets using the largest LLMs, such as the Chinese SMILE (Qiu et al., 2024) and Korean SoulChat (Chen et al., 2023) datasets. The respective works demonstrate that synthetic datasets can be successfully used for training LLMs to improve in ERG. However, neither study evaluated the general language understanding of the trained models, so it is impossible to ascertain whether the training led to the specialization of the models.
@@
Automated Evaluation of Empathy in Dialogue
Evaluation of empathy in dialogue has primarily focused on human evaluation and lexical overlap such as BLEU (Papineni et al., 2002) or ROUGE (Lin, 2004) or semantic similarity with reference dialogues such as BERTscore Zhang et al. (2019). The former is widely used. However, it is limited by the subjective nature of the task, biases introduced by the evaluation design, and, usually, small sample sizes. Given the highly fluent nature of generative LLM responses, we have found the latter to have little relevance for measuring empathy, in agreement with the findings of Liu et al., 2016. More recent methods rely on model-based empathy metrics, which use classification models trained specifically for measuring empathy in multi-turn conversations such as diff-Epitome (Lee et al., 2022b). diff-Epitome is an evolution of the Epitome classifiers (Sharma et al., 2020), which measure empathy in dialogue on a scale (0-2) from none (0), through weak (1), to strong (2) on three dimensions: empathetic responses (ER), explanations (EX), and interpretations (IP). diff-Epitome uses a similar but continuous scale on the same dimensions by averaging the scores across the entire dataset, thus providing a measure of difference from the ground truth.
In contrast to representation-based evaluation models, Lee et al. (2024) employ feature-based methods for evaluation. These include specificity based on a normalized variant of inverse document frequency (NIDF, See et al. (2019)), which shows emotional connections and language that conveys an understanding of the other’s feelings and experiences Truax and Carkruff (1964). It further includes word choice metrics derived from the NRC Emotion Intensity Lexicon Mohammad (2018), which are intended to capture the consistency of the response with the dialog context; thus, valence, arousal, and intensity are measured for both the prompt and response and the ideal is for them to score similarly. Finally, response diversity is measured using response-tries as consideration of the common pitfall of generation systems to resort to generic responses like “I am so sorry to hear that” Lee et al. (2024). The work also offers an averaging schema to account for the dialog context.
Reflectivity has also been used as a proxy measure for empathy: Lee et al. (2024) used PAIR Min et al. (2022) to measure the reflection level in empathetic utterances.
Human Evaluation of Empathy in Dialogue
Human judges, usually crowd workers or students, have been extensively used to evaluate empathy in dialogue. The judges are instructed to evaluate the generated responses in comparative A/B tests or one model at a time in a scale-based evaluation. For the former, an alternative yes/tie/no rating is common, and the result is derived by majority voting. The latter employs three or more dimensions, such as empathy, relevance, informativeness, coherence, or fluency. The quality of the surveys varies as some studies apply one question per dimension, and the exact description of the physical surveys is rare. The sample size is usually 100-150 samples per model. Some stratify the sample on the grounding emotions such as MIME (Majumder et al., 2020b). We present an overview of the evaluation methods used with EmpatheticDialogues in Table 1.
Study Type Dimensions
EmpatheticDialogues (Rashkin et al., 2019) A/B, 5p Likert empathy, relevance, fluency
MoEL (Lin et al., 2019)
MIME (Majumder et al., 2020b)
EmpDG (Li et al., 2020)
EmoCause/GEE (Kim et al., 2021) A/B, 4p Likert
ESConv (Liu et al., 2021) Likert fluency, identification,
comforting, suggestion
EmpBot (Zaranis et al., 2021) 5p Likert fluency & relevance,
empathy-sentiment,
empathy-emotion
CARE (Wang et al., 2022) A/B, 5p Likert coherence, empathy, informativeness
DiffusEmp (Bi et al., 2023)
CEM (Sahand Sabour, 2021) aspect-wise A/B
EPITOME (Lee et al., 2022b) 3p Likert emotional reactions, interpretations,
explorations
SoulChat (Chen et al., 2023) content, empathy, helpfulness, safety
Table 1: Overview of human evaluation methods in recent related work.
Human evaluation offers valuable gold-standard feedback but not without drawbacks, including subjectivity, limited sample size, and the challenge of designing bias-free surveys. LLM-as-a-judge methods have emerged to address these issues. However, their validity remains contested because surveys such as Bavaresco et al. (2024) point out that these methods introduce several new biases, including the position bias (Koo et al., 2024), form-over-substance (Wu and Aji, 2023), model-relatedness (Baris Schlicht et al., 2024), and length sensitivity (Chen et al., 2024).
Language Understanding Evaluation
An invaluable trait of LLMs is their ability to generalize across various domains, tasks, and languages. Language understanding evaluation compares and ranks different models based on this ability. From the many existing benchmarks, two types are especially relevant for dialog-capable models: live A/B testing using either human judgments such as Chat Arena (Zheng et al., 2023) or using LLM-as-a-judge such as in AlpacaEval (Dubois et al., 2024) and static language understanding benchmarks such the MMLU or the Open LLM leaderboard tasks.
Static benchmarks offer consistency and larger sample sizes. MMLU assesses how instruction-following LLMs learn and apply knowledge across various domains by measuring their performance on 16K different multiple-choice style question tasks from STEM, humanities, and world knowledge. It shows the best correlation with human judgments from the Chat Arena among large static benchmarks Dubois et al. (2024). The Open LLM leaderboard Beeching et al. (2023a) is a popular task aggregator to evaluate and rank LLMs. It comprises six tasks that cover the ability to follow instructions, math, science, language understanding (e.g., sarcasm detection, name disambiguation), world knowledge, and multistep reasoning. See Appendix B for a full description of the tasks.
@@ 3 Experimental Setup
We conduct our experiments using the Zephyr-7B (Tunstall et al., 2023) model, a variant of the Mistral-7B foundational LLM (Jiang et al., 2023) fine-tuned for multi-turn dialogues using the UltraChat dataset (Ding et al., 2023). Beeching et al. (2023a) shows that the smallest variant with 7 billion parameters already provides good general language understanding ability as it matches LLama-2-70B on many benchmarks. Using our preference dataset, we compare the untrained baseline Zephyr model to the model trained with DPO. We explore the hyperparameter configuration space for all training steps to minimize the impact on generalization while improving empathy. For control, we added two other sources of responses: the ground-truth responses from ED and the baseline Zephyr model, where we limited the generated length to be similar to the ground-truth. This allows us to offset responders’ cognitive bias towards longer responses being considered of higher subjective quality (Santhanam et al., 2020; Park et al., 2024).
We evaluate all automated metrics using the pre-defined test set of EmpatheticDialogues of 2,540 dialogues.
@@
Tunstall et al. (2023) showed that SFT is a necessary first step in alignment, ensuring the preference dataset is in-domain for the aligned model. We perform SFT using the standard causal auto-regressive objective using the individual dialogues from EmpatheticDialogues. As per the definition of the dataset, the odd turns are considered the "user prompts" and the even turns the "responses". We limit the SFT training to the even-indexed turns by masking the odd turns to ignore them in the loss-function computation. For computational efficiency, we fine-tune using LoRA adapters (Hu et al., 2021).
@@
For constructing the completion pairs, we leverage a property of the ED dataset: each dialogue is associated with an emotion label. These were used as grounding during the ideation of each dialogue. For the preferred completion, we use the ground truth – the original response. For its rejected counterpart, we use Plutchik’s wheel of emotions (Figure 7) and the derivative emotional dyads (Figure 8, Plutchik, 2001) to find the polar opposite emotion labels resulting in a lookup table (see Appx. A). For each completion, we randomly select one from the group of completions labeled with the opposite label. Because the training stability can suffer from random selection, we draw a fresh random completion for each new training epoch and search for the hyperparameter configuration to offset the repetition of the preferred completions.
Model $\mathbf{\overline{x}}($diff-Epitome)$\downarrow$) ↓ diff-ER $\downarrow$ diff-EX $\downarrow$ diff-IP $\downarrow$ FBert $\uparrow$
@@
[…]
Table 2: Model-based empathy metrics diff-Epitome and semantic similarity metric BERTscore with EmpatheticDialogues. All presented results are means ($\pm$ SD) of scores over multiple (4) training runs with the same hyperparameters but different seeds (and different random completion draw for DPO). Bold indicates best results and significant difference ($p<$1e-5) to the next best result using permutation tests.
Model MMLU (5s) $\uparrow$ Open LLM (Acc) $\uparrow$
baseline: Zephyr-7B-sft-full .$588\pm.00$$276\pm.00$
Zephyr-7B + SFT .$585\pm.00$$270\pm.00$
Zephyr-7B + SFT + DPO .$584\pm.00$$266\pm.00$
Table 3: Language understanding metrics: MMLU and Open LLM Leaderboard accuracy over all its tasks. All presented results are means ($\pm$ SD) of scores over multiple (4) training runs with the same hyperparameters but different seeds (and different random completion draw for DPO). With the SFT (lr=1-e5, $\alpha$=64, epoch=1, batch_size=64, rank=1024) and DPO (lr=1e-6, $\beta$=0.05, epoch=3) hyperparameters, we didn’t find a significant difference using the McNemar’s test and SD in all cases was ($<0.001$).
Automated Empathy Evaluation
For measuring empathy, we use the diff-Epitome metric, which measures the difference between levels of empathy of the ground truth and generations. We consider a lower difference from the ground truth as better. We used diff-Epitome also for model selection during the fine-tuning using a validation sample of the training data.
In addition, we consider the set of feature-based metrics from Lee et al. (2024) for further analysis on the generations for the test set, which include NIDF for specificity, lexicon-based valence-arousal-intensity scores (VAD), and diversity metric based on response-tries. The VAD metrics measure the difference between the prompt’s and response’s scores, smaller differences being the desired outcome. To reflect this, we consider the score distance, i.e., the absolute value of the difference, before aggregating the results per prompt-response pair (whereas the prior work aggregated the prompt score minus the response score).
We excluded common lexical-overlap and vector-similarity metrics because our preliminary experiments found them uncorrelated with the perceived quality of the generated responses. It was especially evident with the largest LLMs, such as GPT4 (Achiam et al., 2023) and Claude (Anthropic, 2023), which generated high-quality responses yet received low scores from overlap metrics. However, we keep the semantic similarity BERTscore as a sanity check and for increased metric diversity. Furthermore, we could not reproduce the PAIR model, which is included among Lee et al. (2024)’s metrics.
@@
We measure the general language understanding of the models with MMLU and tasks from the Open LLM leaderboard using the implementation in lm-evaluation-harness Gao et al. (2021).
Human Evaluation
Taking into account critique of human evaluation (Lahnala et al., 2022; Lee et al., 2024), and building off of the lack of specificity described in Coll et al. (2017), we adapted the existing EmpatheticDialogues singular question about situational empathy into four dimensions: emotion understanding, situational appropriateness, response naturalness, and conversational engagingness. Non-empathy associated metrics were combined into a single question of understandability and consistency. The evaluators were hired using the Prolific platform Prolific (2024) and screened for education in Psychology or Social Work.
The evaluations of the four empathy dimensions were performed on a continuous scale from 0-100 where evaluators were asked to rate their degree of agreement to Likert-like statements for each dimension (Ji et al., 2022). Understandability and consistency was evaluated with a checkbox. The evaluation sample comprises 135 examples, each of the 32 emotions in ED represented at least four times. Each example was evaluated by three annotators in batches of 15 examples which were randomized to contain blocks of outputs from different models. In total 108 batches (1620 samples) were evaluated. Sample averaged annotator ratings for each dimension and model, which are averaged, form the final score. We used the pairwise t-test and Cohen’s d on $\alpha=0.05$ to test the difference between different response sources. See Appendix D for more details on the experiment design.